# Cavity-Altered Superconductivity

### v1

@@ Methods
Heterostructures were fabricated by placing vdW materials, on the order of 10 nm to 120 nm thick, on large flat surfaces of -ET single crystals (ac crystallographic plane, Supplementary Information Section 2). Measurements were performed on five devices (Supplementary Information Section 3). Data in Fig 1 is from device 1, Fig 2 from device 2, Fig 3(b-d) from device 1, and Fig 3a from device 5. Fig 4 is a compilation of results from devices 1, 3, and 4.
@@
hBN/BSCCO samples were prepared by exfoliating and assembling BSCCO and hBN ctystals in a golvebox filled with inert gas ( ppm; ppm). BSCCO was first exfoliated on SiO2 substrates. hBN flakes were then exfoliated on a PDMS stamp and transferred onto BSCCO at room temperature.
@@
The nano-infrared scattering experiments were performed using a home-built cryogenic scattering-type scanning near-field optical microscope (s-SNOM) housed in an ultra-high vacuum chamber with a base pressure of torr. The s-SNOM is based on a tapping-mode atomic force microscope (AFM) with the tapping frequency and amplitude of the AFM set to approximately 285 kHz and 70 nm, respectively. The s-SNOM operates by scattering tightly focused laser light from a sharp AFM tip, enabling near-field interactions to be probed with nanometer-scale resolution. The laser source is a tunable quantum cascade laser from Daylight Solutions. The laser beam was focused onto the metallized AFM tip using a parabolic mirror with a 12-mm focal length. The back-scattered light was registered by a mercury cadmium telluride detector and demodulated following a pseudoheterodyne scheme. The signal was demodulated at the th harmonic of the tip tapping frequency, yielding background-free near-field signal. In this work, we selected =3 and 4 to eliminate the far-field background.
@@
MFM measurements were performed in an Attocube cantilever-based cryogenic AFM (attoAFM I with attoLIQUID 2000 cryostat), where the microscope sits in a helium exchange gas at low temperature. Nanosensors PPP-MFMR probes with hard magnetic coating and force constants near N/m were used for all measurements. Measurements were taken on different samples using different tips, all with resonance frequencies ranging from 75.5 kHz to 79.4 kHz. Measurements recorded the resonant frequency shift of the probe cantilever () and were converted to via [41], using the force constants provided by the manufacturer for each cantilever and the measured resonant frequency far from the sample surface, . The tip oscillation amplitude was typically 50 nm - 70 nm. hBN is non-magnetic and therefore does not impact the detection of the Meissner effect over hBN/-ET. RuCl3 is antiferromagnetic at low temperature, but has a low susceptibility.
To minimize noise and instrument drift, was typically averaged over multiple measurements repeated at each point and taken with high resolution (typically 1-2 nm to define the surface accurately). During the measurements, the tip was electrically biased to compensate for the local contact potential difference and thus minimize electrostatic forces. The average curves were then smoothed with a uniform filter with width 30 to 50 pixels to reduce the appearance of noise. For the -ET heterostructures, a linear background was subtracted by fitting the data far from the sample surface where the Meissner force is negligible. For the BSCCO sample, a background measured on the silicon substrate was subtracted (Supplementary Information Section 5). For proper comparison in Fig 1, we have compensated for the thickness of the hBN and RuCl3 such that the measurements for hBN/-ET, bare -ET, and RuCl3/-ET all have the -ET surface at zero.
Constant height MFM images were taken with the tip at a fixed height above the plane of the -ET surface. The images shown are raw data, except for the linear conversion from to and an overall constant offset to calculate . During the measurements, the tip was held at a fixed electric bias to compensate for the contact potential difference at one point in the field of view.
The data points in Fig. 4 were obtained either by fitting curves or by converting constant height MFM images to maps of the local , as described in Supplementary Information Section 4.

### v3

@@ Methods
Heterostructures were fabricated by placing vdW materials, on the order of 10 nm to 120 nm thick, on large flat surfaces of -ET single crystals (ac crystallographic plane, Supplementary Information Section 2). Measurements were performed on five devices (Supplementary Information Section 3). Data in Fig. 1 are from device 1, in Fig. 2 from device 2, in Fig. 3(b-d) from device 1, and in Fig. 3a from device 5. Fig. 4 is a compilation of results from devices 1, 3, and 4.
@@
hBN/BSCCO samples were prepared by exfoliating and assembling BSCCO and hBN crystals in a glovebox filled with inert gas ( ppm; ppm). BSCCO was first exfoliated on SiO2 substrates. hBN flakes were then exfoliated on a PDMS stamp and transferred onto BSCCO at room temperature.
@@
The nano-infrared scattering experiments were performed using a home-built cryogenic scattering-type scanning near-field optical microscope (s-SNOM) housed in an ultra-high vacuum chamber with a base pressure of torr. The s-SNOM is based on a tapping-mode atomic force microscope (AFM) with the tapping frequency and amplitude of the AFM set to approximately 285 kHz and 70 nm, respectively. The s-SNOM operates by scattering tightly focused laser light from a sharp AFM tip, enabling near-field interactions to be probed with nanometer-scale resolution. The laser source is a tunable quantum cascade laser from Daylight Solutions. The laser beam was focused onto the metallized AFM tip using a parabolic mirror with a 12-mm focal length. The backscattered light was registered by a mercury cadmium telluride detector and demodulated following a pseudoheterodyne scheme. The signal was demodulated at the th harmonic of the tip tapping frequency, yielding a background-free near-field signal. In this work, we selected =3 and 4 to eliminate the far-field background.
@@
MFM measurements were performed in an Attocube cantilever-based cryogenic AFM (attoAFM I with attoLIQUID 2000 cryostat), where the microscope sits in a helium exchange gas at low temperature. Nanosensors PPP-MFMR probes with hard magnetic coatings and force constants near N/m were used for all measurements. Measurements were taken on different samples using different tips, all with resonance frequencies ranging from 75.5 kHz to 79.4 kHz. Measurements recorded the resonant frequency shift of the probe cantilever () and were converted to via [40], using the force constant provided by the manufacturer for each cantilever and the measured resonant frequency far from the sample surface, . The tip oscillation amplitude was typically 50 nm - 70 nm. hBN is non-magnetic and therefore has no impact on the detection of the Meissner effect over hBN/-ET. RuCl3 is antiferromagnetic at low temperature, but has a low susceptibility.
To minimize noise and instrument drift, was typically averaged over multiple measurements repeated at each point and taken with high resolution (typically 1-2 nm to define the surface accurately). During the measurements, the tip was electrically biased to compensate for the local contact potential difference and thus minimize electrostatic forces. The average curves were then smoothed with a uniform filter with width 30 to 50 pixels to reduce the appearance of noise. For the -ET heterostructures, a linear background was subtracted by fitting the data far from the sample surface where the Meissner force is negligible. For the BSCCO sample, a background measured on the silicon substrate was subtracted (Supplementary Information Section 5). For proper comparison in Fig. 1, we have compensated for the thickness of the hBN and RuCl3 microcrystals such that the measurements for hBN/-ET, bare -ET, and RuCl3/-ET all have the -ET surface at zero.
Constant-height MFM images were taken with the tip at a fixed height above the plane of the -ET surface. The images shown are raw data, except for the linear conversion from to and an overall constant offset to calculate . During the measurements, the tip was held at a fixed electric bias to compensate for the contact potential difference at one point in the field of view.
The data points in Fig. 4 were obtained either by fitting curves or by converting constant-height MFM images to maps of the local , the decay length for currents running parallel to the surface. was converted to superfluid density via the proportionality to , as described in Supplementary Information Section 4.
@@
Data availability
All datasets generated and analyzed during this study are available upon request from the authors.
Acknowledgments We acknowledge helpful discussions with O. Auslaender. This work was supported by the European Research Council (ERC-2024-SyG- 101167294 ; UnMySt), the Cluster of Excellence Advanced Imaging of Matter (AIM), Grupos Consolidados (IT1249-19) and SFB925, We acknowledge support from the Max Planck-New York City Center for Non-Equilibrium Quantum Phenomena. The Flatiron Institute is a division of the Simons Foundation. Support for hexagonal boron nitride crystal growth was provided by the Office of Naval Research, Award No. N00014-22-1-2582. K.K. acknowledges the support of the JSPJ KANENHI (21K18144), the JSPS Core-to-Core Program (JPJSCCA20240001) and the Alexander von Humboldt Foundation. The work at BNL was supported by the US Department of Energy, office of Basic Energy Sciences, contract no. DOE-SC0012704. D.M. and M.C. acknowledge support from the Gordon and Betty Moore Foundation’s EPiQS Initiative, Grant GBMF9069. D.M.K. and M.A.S. acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - 531215165 (Research Unit ‘OPTIMAL’). M.A.S. was funded by the European Union (ERC, CAVMAT, project no. 101124492). T.O. acknowledges support from the JSPS Overseas Research Fellowships. D.S. was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00333664). J. Y. and Q.L were supported by the Office of Basic Energy Sciences, Materials Sciences and Engineering Division, U.S. Department of Energy (DOE) under Contract No. DE-SC0012704. We acknowledge project/application support by the Max Planck Computing and Data Facility.
Author contributions I.K., T.A.W., J.X., and G.P. collected and analyzed MFM data. I.K. and T.A.W. wrote the manuscript. S.Z. and S.S.Z. collected SNOM measurements. S.Z and J.Z. analyzed SNOM data. S.Z. simulated the HPhP dispersion for hBN on -ET. D.S. and B.S.Y.K. fabricated the heterostructures with input from C.R.D. G.P. performed SEM imaging of the MFM tips. J.Y. and Q.L. performed SQUID magnetometry measurements. T.O. provided simulation of the electric field in a z-dependent superfluid density scenario. J.H.E. grew the hBN. M.J., R.P., and S.W. provided discussion and interpretation of the data, as well as formulation of the idea. K.M. and K.K. grew the -ET crystals. G.G grew the BSCCO crystal. D.M. and M.C. grew the RuCl3 crystal. M.B. and A.C. provided discussion and interpretation of the data. D.M.K. provided guidance on fitting and participated in the discussion on mechanisms. D.S. performed the ab initio calculations under the supervision of A.R. D.S., M.H.M., and A.R. analyzed the ab initio results and developed a microscopic model for the quenching of superconductivity. M.A.S. provided theoretical input on the interpretation of results and contributed to the manuscript. A.N.P. advised on measurements and data interpretation. D.N.B. advised and supervised the project. All authors discussed the results and contributed to the final paper.
Competing interests The authors declare no competing interests.
Correspondence and requests for materials should be addressed to Itai Keren, Tatiana Webb, Shuai Zhang or Dmitri Basov.